• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name jroche

    Log entry time 18:15:39 on September20,2010

    Entry number 332627

    keyword=ADC need to be retimed.

    The plots below show the ADC spectra for some of the S2 scintillators for run 6334 (plot 2) taken on September 14 and for run 6568 (plot 1) taken today. Both run are taken with ps1 triggers. Looks to me that the ADC gate is now missing the signals.
    In between September and today, Hisham recabled the triggers on September 16 and from what I understand he is waiting for the calo trigger to finish his work. If I understand these two plots correctly, retiming the ADC gate will have to be done when final timing is frozen.

    FIGURE 1

    FIGURE 2